help
Solve the quadratic equation by completing the square.
x^2 +8x+11=0


Found 3 solutions by ewatrrr, ikleyn, josgarithmetic:
Answer by ewatrrr(24785)

 
Hi,
Solve the quadratic equation by completing the square.
x^2 +8x+11 + 5 = 5
x^2 +8x+16 = 5
(x+4)^2 = 5
(x+4) = ± √5
x = -4 ± √5
checking our work with Quadratic Formula
x =(-8 ± √(64- 44))/2
x =( -8 ± √(20))/2
x =( -8 ± 2√(5))/2
x = -4 ± √(5)
Wish You the Best in your Studies.

Answer by ikleyn(52780)

.

On completing the square,  see the lesson
    - HOW TO solve quadratic equation by completing the square - Learning by examples
in this site.

Learn the subject from there.



Answer by josgarithmetic(39617)